OpenStudy (anonymous):

If f(x) = 4x – 1 and g(x) = x^2 + 2x + 2, what is f(g(x))?

OpenStudy (anonymous):

What have you tried? :)

OpenStudy (anonymous):

I tried F.O.I.L.ing 4x-1(x^2+2x+2) .

OpenStudy (anonymous):

See if this helps: if \( f(x) = x +2\) and \(g(x)= x^2\) then \(f(g(x)) = x^2+2\)

OpenStudy (anonymous):

In your case, it will be something like \( 4(x^2 + 2x + 2)+1 \) now simplify :)

OpenStudy (anonymous):

I don't get it. wouldn't it be x^3+2?

OpenStudy (anonymous):

No, FOIL has not thing to do with composition.

OpenStudy (anonymous):

ohhh I didn't know that.

OpenStudy (anonymous):

So is the answer 4x^2+8x+9?

OpenStudy (anonymous):

Yes, but I made a lil mistake in sign 4(x^2 + 2x + 2)-1 so that's 4x^2+8x+7

OpenStudy (anonymous):

Ohh alright thank you :D

OpenStudy (anonymous):

Could you help me with another problem?

OpenStudy (anonymous):

if f(x)=x-2/2, then what is f^-1(x)?

OpenStudy (anonymous):

Do you know what is inverse function?

OpenStudy (anonymous):

and are you sure it's 2/2 ?

OpenStudy (anonymous):

Not really... & yeah it's x-2 over 2.

OpenStudy (anonymous):

2x+2

OpenStudy (anonymous):

But you should read about inverse function in detail. There are some subtle things here.

OpenStudy (anonymous):

Thank you & will do :)

OpenStudy (anonymous):

Glad to help :)